# HoTPP Benchmark: Are We Good at the Long Horizon Events Forecasting?

### v3

@@ 4.1. Datasets
We for the first time combine datasets from a wide range of domains such as financial transactions, social networks, and medical records into a single evaluation benchmark. Specifically, we provide evaluation results on a transactional dataset (Babaev et al., 2022), the MIMIC-IV medical dataset (Johnson et al., 2023), and social network datasets (Retweet (Zhao et al., 2015), StackOverflow (Jure, 2014), and Amazon (Jianmo, 2018)). These datasets represent diverse underlying processes: social network data is influenced by cascades (Zhao et al., 2015), medical records exhibit repetitive patterns and transactional data reflects daily activities, combining regularity with significant uncertainty.
@@
The dataset statistics are presented in Table 1. The Transactions dataset has the longest average sequence length and the largest number of classes, while MIMIC-IV contains the highest number of sequences. Retweet is medium size, whereas Amazon and StackOverflow can be considered small datasets. Further datasets details are reported in Appendix C.4 and Appendix F.
4.2. Methods
@@
Intensity-free approaches. We implement the IFTPP method, which combines mean absolute error (MAE) of the time step prediction with cross-entropy categorical loss for labels (Shchur et al., 2019; Padhi et al., 2021; McDermott et al., 2024). IFTPP uses GRU reccurrent network as a backbone (Cho et al., 2014). We also evaluate IFTPP-T model with a transformer backbone (Radford et al., 2019).
Intensity-based approaches. We implement RMTPP (Du et al., 2016) as an example of the TPP approach with a traditional RNN. We add NHP (Mei and Eisner, 2017), based on a continuous time LSTM architecture. We also evaluate the AttNHP approach that utilizes a continuous time transformer model (Yang et al., 2022) and ODE (Rubanova et al., 2019), which applies ordinary differential equations for temporal dynamics modeling.
Reranking. We implement HYPRO (Xue et al., 2022), which generates multiple hypotheses with RMTPP and selects the best sequence using a contrastive approach.
Diffusion. We include the Diffusion approach (Zhou et al., 2025), which extends previous research in language modeling (Li et al., 2022). Diffusion model iteratively refines predicted sequence, simultaneously generating all events within the specified horizon. Notably, HoTPP provides the first public implementation of this approach.
We also explored ContiFormer (Chen et al., 2023) and NJDTPP (Zhang et al., 2024), but excluded them from our final comparison for the reasons detailed in Appendix A.1.
Additional details on implemented methods are provided in Appendix A, with training specifics outlined in Appendix D and hyperparameters listed in Appendix E.
5. Experiments
@@
5.1. Long-horizon Prediction Quality
@@
Figure 5. Comparison of predicted label entropies, with the dashed horizontal line indicating the ground truth entropy.
5.2. Metrics Comparison
@@
Table 3. Examples of predicted sequences (labels only).
Method Transactions MIMIC-IV Retweet Amazon StackOverflow IFTPP 6, 6, 6, 6, 6, 6 14, 12, 4, 6, 11, 10, 1, 3, 5, 8, 9, 16 0, 0, 0, 0, 0, 0, 0, 0, 0, 0 0, 11, 0, 11, 0, 0, 0, 0, 0, 0, 0, 0 7, 17, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3 RMTPP 6, 6, 6, 6, 6, 6 14, 4, 6, 11 0, 0, 0, 0, 0, 0, 0 11, 11, 11, 11, 11, 11, 11, 11, 11, 11, 11, 11 17, 3, 3, 3, 3, 3 NHP 6, 6, 6, 6, 6 4, 6 1, 1, 1, 1, 1, 1, 1, 1 5, 5, 5, 5, 5, 5, 5, 5, 5, 5 3, 3, 3, 3, 3, 3 AttNHP 1, 1, 1, 1, 1 4, 6, 10, 1 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 3, 0, 4, 1, 1, 1, 0, 0, 0, 0, 0, 0 3, 3, 3, 3, 3, 3, 3 ODE 1, 6, 6, 6, 6, 6 4, 6, 1, 3, 5, 2, 7, 4, 6 0, 0, 0, 0, 0, 0, 0 5, 5, 5, 5, 5, 5, 5, 5, 5, 5 3, 3, 3, 3, 3, 3 HYPRO 3, 6, 6, 1, 6, 1, 6 24, 1, 3, 5, 17, 8, 2, 20 0, 0, 1, 0, 0, 1, 0 1, 11, 0, 0, 11, 11, 11, 5, 1, 0, 5, 11 17, 5, 3, 5, 13, 3, 5 Diffusion 6, 6, 6, 6, 6, 6, 6 14, 4, 6, 1, 1, 3, 3, 3, 8, 3, 2, 2 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 0, 0, 8, 5, 11, 11, 5, 5, 0, 5, 5, 11 3, 3, 3, 3, 3, 3, 3, 3
5.3. Mode Collapse
We found that most models tend to predict a single label or repetitive patterns in long-horizon tasks, as shown in Table 3. To quantify this effect, we measured the entropies of predicted labels (Fig.5). Autoregressive methods generally exhibit low prediction diversity across all datasets except MIMIC-IV, where strong dependencies enable the model to learn typical sequences. Meanwhile, HYPRO, which samples from RMTPP with an increased temperature, substantially enhances label diversity. Diffusion similarly boosts diversity except on the Transactions dataset. Overall, most methods, except HYPRO, experience mode collapse in long-horizon prediction, although horizon-based approaches partly mitigate this issue. Sampling with an increased temperature can potentially solve this issue by offering a trade-off between predictions quality and diversity. Additional examples of predicted sequences are presented in Appendix G.
6. Limitations and Future Work
The proposed benchmark has several shortcomings. The training process for some methods is relatively slow. For example, a continuous time LSTM from the NHP method lacks an effective open-source GPU implementation, which could be developed in the future. Autoregression can also be optimized using specialized GPU implementations, which would significantly impact HYPRO training. Additionally, the list of implemented methods can be extended, for example, by incorporating multi-task models (Deshpande et al., 2021).
@@ 6. Limitations and Future Work
On the other hand, the T-mAP metric can be applied in domains beyond event sequences, such as action recognition (Su and Hua, 2017). T-mAP can potentially offer better timestamp evaluation in these domains than evaluating indices with OPW. T-mAP can also provide a more natural hyperparameter selection regarding the modeling horizon and maximum allowed time error. Our theoretical justification of T-mAP can potentially be adapted for mAP estimation algorithms in computer vision.
7. Conclusion
In this paper, we propose the HoTPP benchmark to assess the quality of long-horizon events forecasting. HoTPP provides optimized implementations of the state-of-the-art methods ranging from statistical baselines to ODE and Diffusion (Sec.4), along with diverse metrics for both next-event and long-horizon prediction (Sec.3). Our extensive evaluations using diverse datasets and various predictive models reveal a critical insight: high accuracy in next-event prediction typically does not correlate with superior performance in horizon prediction (Sec. 5.1). Moreover, most existing methods suffer from mode collapse when forecasting longer sequences (Sec. 5.3). These findings underscore the need for developing specialized models. By emphasizing long-term predictive capabilities, HoTPP aims to stimulate the development of more robust and reliable event-sequence models. This advance has the potential to substantially enhance practical applications in various industries, driving both innovation and more informed decision-making.
@@ Appendix A Modeling Marked Temporal Point Processes
The derivation is provided in (Rizoiu et al., 2017). Different TPPs are characterized by their intensity function $\lambda(t_{i})$. In Poisson and non-homogeneous Poisson processes, the intensity function is independent of previous events, meaning event occurrences depend solely on external factors. In contrast, self-exciting processes are characterized by previous events increasing the intensity of future events. A notable example of a self-exciting process is the Hawkes process, in which each event linearly affects the future intensity:
(5) $\lambda(t_{i})=\lambda_{0}(t_{i})+\sum\limits_{k=1}^{i-1}\phi(t_{i}-t_{k}),$
@@
where $\lambda_{0}(t_{i})\geq 0$ is the base intensity function, independent of previous events, and $\phi(x)\geq 0$ is the so-called memory kernel function. Given the intensity function $\lambda(t_{i})$, predictions are typically made by sampling or expectation estimation. Sampling is usually performed using the thinning algorithm, a specific rejection sampling approach. For details on the implementation of sampling, please refer to (Rizoiu et al., 2017).
@@
Recent research has focused on modeling complex intensity functions. Various neural network architectures have been adapted to address this problem. These approaches differ in the type of neural network used and the model of the intensity function. Neural architectures range from simple RNNs and Transformers to specially designed continuous-time models like NHP (Mei and Eisner, 2017) and Neural ODEs (Rubanova et al., 2019). The intensity function between events can be modeled as a sum of intensities from previous events, as in Hawkes processes, or by directly predicting the inter-event intensity given the context embedding.
@@
Intensity-free modeling. Some methods evaluate the next event time distribution without using intensity functions. For example, intensity-free (Shchur et al., 2019) represents the distribution as a mixture of Gaussians or through normalizing flows. Instead of predicting the distribution directly, other approaches solve a regression problem using MAE or RMSE loss. However, it can be shown that both MAE and RMSE losses are closely related to distribution prediction. Specifically, RMSE is analogous to log-likelihood optimization with a Normal distribution, and MAE is similar to the log-likelihood of a Laplace distribution (Bishop, 1994). In our experiments, we evaluated a model trained with MAE loss as an example of an intensity-free method.
@@
Rescoring with HYPRO. HYPRO (Xue et al., 2022) is an extension applicable to any sequence prediction method capable of sampling. HYPRO takes a pretrained generative model and trains an additional scoring module to select the best sequence from a sample. It is trained with a contrastive loss to distinguish between the generated sequence and the ground truth. HYPRO generates multiple sequences with a background model during inference and selects the best one by maximizing the estimated score. Although HYPRO is intended to improve quality compared to simple sampling, it is unclear whether HYPRO outperforms expectation-based prediction. In our work, we apply HYPRO to the outputs of the RMTPP model.
@@
Diffusion. Recently, diffusion-based approaches have been applied to NLP tasks (Li et al., 2022). This idea has been extended to event sequences by adapting best practices from NLP (Zhou et al., 2025). In such methods, events are first mapped to continuous embeddings, where the diffusion process operates, then mapped back to event space. The process begins with embeddings sampled from Gaussian noise and iteratively refines them using historical event embeddings. Finally, the refined embeddings are projected to event times and labels through a softmax-based layer. Training involves a combination of MSE loss in the continuous space, MAE loss for time reconstruction, and cross-entropy loss for label prediction. In HoTPP, we provide the first public implementation of the approach proposed by (Zhou et al., 2025).
@@
A.1. Methods Excluded from Comparison
We attempted to incorporate ContiFormer (Chen et al., 2023) into our benchmark but encountered multiple challenges. The official implementation provides scripts only for training on a toy spiral dataset, making it impossible to reproduce the method’s performance on real-world data. Moreover, training even on this toy example required approximately three hours on an NVIDIA RTX 4060, suggesting that ContiFormer cannot be efficiently scaled to large datasets such as Transactions. In addition, contrary to the claims in the paper, ODE (Rubanova et al., 2019) and continuous-time transformers (Yang et al., 2022) were already discussed in earlier work. Therefore, we excluded ContiFormer from our benchmark in favor of ODE and AttNHP.
We also attempted to include NJDTPP (Zhang et al., 2024) in our benchmark, but encountered several challenges. Even on StackOverflow, our smallest dataset, the inference time spanned two hours for a single test set prediction. Moreover, after running the official implementation on StackOverflow, we observed accuracy values that were inconsistent with the paper and fell below most baseline results. We also identified data leakage during inference, where future event timestamps were used to predict the subsequent event. Consequently, we chose not to include NJDTPP in our comparisons.
@@ Appendix B T-mAP Computation and Proofs
(6) $\forall h\forall m\in M(\mathcal{G})\exists m_{h}\in M(\mathcal{G}_{h}):m_{h}\subset m.$
@@
Consider the case when matching $m(\mathcal{G})$ contains an edge $(i,j)$. Let $\hat{m}_{h}=m(\mathcal{G})\setminus\{(i,j)\}$, i.e., the matching without edge $(i,j)$. If it is optimal for the graph $\mathcal{G}_{h}$, then it will satisfy the theorem. Otherwise, there is an optimal matching $m_{h}$ with either (a) more vertices or (b) a smaller total weight than $\hat{m}_{h}$.
@@
(7) $\mathrm{T\text{-}mAP}(\{\mathcal{S}_{p}^{i}\},\{\mathcal{S}_{gt}^{i}\})=\frac{1}{L}\sum\limits_{l=1}^{L}\mathrm{AP}(\{\mathcal{S}_{p}^{i}\},\{\mathcal{S}_{gt}^{i,l}\}).$
@@
(8) $\mathrm{AP}=\sum\limits_{i}\left(\mathrm{Rec}_{i}-\mathrm{Rec}_{i-1}\right)\mathrm{Prec}_{i},$
@@
(9) $C=\operatorname*{arg\,min}\limits_{m\in M(\mathcal{G})}\sum\limits_{(i,j)\in m}(-s^{p}_{i}).$
@@
Appendix C HoTPP Benchmark Details
C.1. Architecture
HoTPP incorporates best practices from extensible and reproducible ML pipelines. It leverages PyTorch Lightning (Falcon and The PyTorch Lightning team, 2019) as the core training library, ensuring reproducibility and portability across various computing environments. Additionally, HoTPP utilizes the Hydra configuration library (Yadan, 2019) to enhance extensibility. The overall architecture is illustrated in Fig. 7. HoTPP supports both discrete-time and continuous-time models as well as RNN and Transformer architectures. Implementing a new method requires only adding essential components, while the rest can be configured through Hydra and configuration files.
Figure 7. The architecture of the HoTPP library.
C.2. Metrics
MTPP models are typically evaluated based on the accuracy of next-event predictions. Common metrics include mean absolute error (MAE) or root mean square error (RMSE) for time shifts and accuracy or error rate for label predictions. Some studies also assess test set likelihood as predicted by the model, but we do not include this measure as it is intractable for the IFTPP and Diffusion models. Previous works have advanced long-horizon evaluation using OTD (Mei et al., 2019; Xue et al., 2023). In addition to these metrics, we evaluate the novel T-mAP metric, which addresses the shortcomings of previous metrics, as discussed in Section 3.1.
C.3. Backbones
We use three types of architectures in our experiments. The first is the GRU network (Cho et al., 2014), one of the top-performing neural architectures for event sequences (Babaev et al., 2022). We also implement the continuous-time LSTM (CT-LSTM) from the NHP method (Mei and Eisner, 2017). Since CT-LSTM requires a specialized loss function and increases training time, we use it exclusively with the NHP method, preferring GRU in other cases. An additional advantage of GRU is that its output is equal to its hidden state, simplifying the estimation of intermediate hidden states for autoregressive inference starting from the middle of a sequence. Lastly, we implement the AttNHP continuous-time transformer model (Yang et al., 2022).
C.4. Datasets

### v4

@@ 4.1 Datasets
We for the first time combine datasets from a wide range of domains such as financial transactions, social networks, and medical records into a single evaluation benchmark. Specifically, we provide evaluation results on a transactional dataset [1], the MIMIC-IV medical dataset [11], and social network datasets (Retweet [39], StackOverflow [13], and Amazon [10]). These datasets represent diverse underlying processes: social network data is influenced by cascades [39], medical records exhibit repetitive patterns and transactional data reflects daily activities, combining regularity with significant uncertainty.
@@
The dataset statistics are presented in Table 1. The Transactions dataset has the longest average sequence length and the largest number of classes, while MIMIC-IV contains the highest number of sequences. Retweet is medium size, whereas Amazon and StackOverflow can be considered small datasets. Further datasets details are reported in F and H.
4.2 Methods
@@
  • Intensity-free approaches. We implement the IFTPP method, which combines mean absolute error (MAE) of the time step prediction with cross-entropy categorical loss for labels [27, 22, 18]. IFTPP uses GRU reccurrent network as a backbone [5]. We also evaluate IFTPP-T model with a transformer backbone [23].
  • Intensity-based approaches. We implement RMTPP [7] as an example of the TPP approach with a traditional RNN. We add NHP [19], based on a continuous time LSTM architecture. We also evaluate the AttNHP approach that utilizes a continuous time transformer model [36] and ODE [25], which applies ordinary differential equations for temporal dynamics modeling.
  • Reranking. We implement HYPRO [34], which generates multiple hypotheses with RMTPP and selects the best sequence using a contrastive approach.
  • Diffusion. We include the Diffusion approach [40], which extends previous research in language modeling [15]. Diffusion model iteratively refines predicted sequence, simultaneously generating all events within the specified horizon. Notably, HoTPP provides the first public implementation of this approach.
We also explored ContiFormer [4] and NJDTPP [38], but excluded them from our final comparison for the reasons detailed in F. Additional details on implemented methods are provided in A, with training specifics outlined in I and hyperparameters listed in J.
Table 2: Evaluation results. The best result is shown in bold. The mean and standard deviation of each metric computed during five runs with different random seeds are reported.
5 Experiments
@@
5.1 Long-horizon Prediction Quality
@@
5.2 Metrics Comparison
@@
5.3 Mode Collapse
Figure 5: Comparison of predicted label entropies for different methods, with the dashed horizontal line indicating the ground truth entropy. The results for all 5 datasets are presented in K.
We found that most models tend to predict a single label or repetitive patterns in long-horizon tasks (see examples from K). To quantify the effect of mode collapse, we measured the entropies of predicted labels, as illustratedin Figure 5. Autoregressive methods generally exhibit low prediction diversity across all datasets except MIMIC-IV, where strong dependencies enable the model to learn typical sequences. Meanwhile, HYPRO, which samples from RMTPP with an increased temperature, substantially enhances label diversity. Diffusion similarly boosts diversity except on the Transactions dataset. Overall, most methods, except HYPRO, experience mode collapse in long-horizon prediction, although horizon-based approaches partly mitigate this issue. Sampling with an increased temperature can potentially solve this issue by offering a trade-off between predictions quality and diversity.
6 Conclusion
In this paper, we proposed the HoTPP benchmark to assess the quality of long-horizon events forecasting. HoTPP provides optimized implementations of the state-of-the-art methods ranging from statistical baselines to ODE and Diffusion (Sec.4), along with diverse metrics for both next-event and long-horizon prediction (Sec.3). Our extensive evaluations using diverse datasets and various predictive models reveal a critical insight: high accuracy in next-event prediction typically does not correlate with superior performance in horizon prediction (Sec. 5.1). Moreover, most existing methods suffer from mode collapse when forecasting longer sequences (Sec. 5.3). These findings underscore the need for developing specialized models. By emphasizing long-term predictive capabilities, HoTPP aims to stimulate the development of more robust and reliable event-sequence models. This advance has the potential to substantially enhance practical applications in various industries, driving both innovation and more informed decision-making.
7 Limitations and Future Work
The proposed benchmark has several shortcomings. The training process for some methods is relatively slow. For example, a continuous time LSTM from the NHP method lacks an effective open-source GPU implementation, which could be developed in the future. Autoregression can also be optimized using specialized GPU implementations, which would significantly impact HYPRO training. Additionally, the list of implemented methods can be extended, for example, by incorporating multi-task models [6].
@@ 7 Limitations and Future Work
On the other hand, the T-mAP metric can be applied in domains beyond event sequences, such as action recognition [28]. T-mAP can potentially offer better timestamp evaluation in these domains than evaluating indices with OPW. T-mAP can also provide a more natural hyperparameter selection regarding the modeling horizon and maximum allowed time error. Our theoretical justification of T-mAP can potentially be adapted for mAP estimation algorithms in computer vision.
@@ Appendix A Modeling Marked Temporal Point Processes
The derivation is provided in [24]. Different TPPs are characterized by their intensity function $\lambda(t_{i})$. In Poisson and non-homogeneous Poisson processes, the intensity function is independent of previous events, meaning event occurrences depend solely on external factors. In contrast, self-exciting processes are characterized by previous events increasing the intensity of future events. A notable example of a self-exciting process is the Hawkes process, in which each event linearly affects the future intensity:
$\lambda(t_{i})=\lambda_{0}(t_{i})+\sum\limits_{k=1}^{i-1}\phi(t_{i}-t_{k}),$ (5)
@@
where $\lambda_{0}(t_{i})\geq 0$ is the base intensity function, independent of previous events, and $\phi(x)\geq 0$ is the so-called memory kernel function. Given the intensity function $\lambda(t_{i})$, predictions are typically made by sampling or expectation estimation. Sampling is usually performed using the thinning algorithm, a specific rejection sampling approach. For details on the implementation of sampling, please refer to [24].
@@
Recent research has focused on modeling complex intensity functions. Various neural network architectures have been adapted to address this problem. These approaches differ in the type of neural network used and the model of the intensity function. Neural architectures range from simple RNNs and Transformers to specially designed continuous-time models like NHP [19] and Neural ODEs [25]. The intensity function between events can be modeled as a sum of intensities from previous events, as in Hawkes processes, or by directly predicting the inter-event intensity given the context embedding.
@@
Intensity-free modeling. Some methods evaluate the next event time distribution without using intensity functions. For example, intensity-free [27] represents the distribution as a mixture of Gaussians or through normalizing flows. Instead of predicting the distribution directly, other approaches solve a regression problem using MAE or RMSE loss. However, it can be shown that both MAE and RMSE losses are closely related to distribution prediction. Specifically, RMSE is analogous to log-likelihood optimization with a Normal distribution, and MAE is similar to the log-likelihood of a Laplace distribution [3]. In our experiments, we evaluated a model trained with MAE loss as an example of an intensity-free method.
@@
Rescoring with HYPRO. HYPRO [34] is an extension applicable to any sequence prediction method capable of sampling. HYPRO takes a pretrained generative model and trains an additional scoring module to select the best sequence from a sample. It is trained with a contrastive loss to distinguish between the generated sequence and the ground truth. HYPRO generates multiple sequences with a background model during inference and selects the best one by maximizing the estimated score. Although HYPRO is intended to improve quality compared to simple sampling, it is unclear whether HYPRO outperforms expectation-based prediction. In our work, we apply HYPRO to the outputs of the RMTPP model.
@@
Diffusion. Recently, diffusion-based approaches have been applied to NLP tasks [15]. This idea has been extended to event sequences by adapting best practices from NLP [40]. In such methods, events are first mapped to continuous embeddings, where the diffusion process operates, then mapped back to event space. The process begins with embeddings sampled from Gaussian noise and iteratively refines them using historical event embeddings. Finally, the refined embeddings are projected to event times and labels through a softmax-based layer. Training involves a combination of MSE loss in the continuous space, MAE loss for time reconstruction, and cross-entropy loss for label prediction. In HoTPP, we provide the first public implementation of the approach proposed by [40].
@@
Methods excluded from comparison. We attempted to incorporate ContiFormer [4] into our benchmark but encountered multiple challenges. The official implementation provides scripts only for training on a toy spiral dataset, making it impossible to reproduce the method’s performance on real-world data. Moreover, training even on this toy example required approximately three hours on an NVIDIA RTX 4060, suggesting that ContiFormer cannot be efficiently scaled to large datasets such as Transactions. In addition, contrary to the claims in the paper, ODE [25] and continuous-time transformers [36] were already discussed in earlier work. Therefore, we excluded ContiFormer from our benchmark in favor of ODE and AttNHP.
We also attempted to include NJDTPP [38] in our benchmark, but encountered several challenges. Even on StackOverflow, our smallest dataset, the inference time spanned two hours for a single test set prediction. Moreover, after running the official implementation on StackOverflow, we observed accuracy values that were inconsistent with the paper and fell below most baseline results. We also identified data leakage during inference, where future event timestamps were used to predict the subsequent event. Consequently, we chose not to include NJDTPP in our comparisons.
@@ Appendix B T-mAP Computation and Proofs
$\forall h\forall m\in M(\mathcal{G})\exists m_{h}\in M(\mathcal{G}_{h}):m_{h}\subset m.$ (6)
@@
[…]
@@
$\mathrm{T\text{-}mAP}(\{\mathcal{S}_{p}^{i}\},\{\mathcal{S}_{gt}^{i}\})=\frac{1}{L}\sum\limits_{l=1}^{L}\mathrm{AP}(\{\mathcal{S}_{p}^{i}\},\{\mathcal{S}_{gt}^{i,l}\}).$ (7)
@@
$\mathrm{AP}=\sum\limits_{i}\left(\mathrm{Rec}_{i}-\mathrm{Rec}_{i-1}\right)\mathrm{Prec}_{i},$ (8)
@@
$C=\operatorname*{arg\,min}\limits_{m\in M(\mathcal{G})}\sum\limits_{(i,j)\in m}(-s^{p}_{i}).$ (9)
@@
Appendix C T-mAP Hyperparameters
The dependency of T-mAP on the $\delta$ parameter for different datasets is presented in Figure 7.
Figure 7: T-mAP dependency on the $\delta$ parameter. The dashed line indicates the selected value.
Appendix D T-mAP for Highly Irregular Sequences
Figure 8: Comparison of simple baselines on the Toy dataset with highly irregular time intervals. MAE and OTD metrics represent error values, while T-mAP measures model quality.
Appendix E T-mAP for Long-tail Prediction Problems
Figure 9: Comparison of various metrics on the Transactions dataset across subsets with fewer event classes.
Appendix F HoTPP Benchmark Details
F.1 Architecture
HoTPP incorporates best practices from extensible and reproducible ML pipelines. It leverages PyTorch Lightning [9] as the core training library, ensuring reproducibility and portability across various computing environments. Additionally, HoTPP utilizes the Hydra configuration library [35] to enhance extensibility. The overall architecture is illustrated in Fig. 10. HoTPP supports both discrete-time and continuous-time models as well as RNN and Transformer architectures. Implementing a new method requires only adding essential components, while the rest can be configured through Hydra and configuration files.
Figure 10: The architecture of the HoTPP library.
F.2 Metrics
MTPP models are typically evaluated based on the accuracy of next-event predictions. Common metrics include mean absolute error (MAE) or root mean square error (RMSE) for time shifts and accuracy or error rate for label predictions. Some studies also assess test set likelihood as predicted by the model, but we do not include this measure as it is intractable for the IFTPP and Diffusion models. Previous works have advanced long-horizon evaluation using OTD [20, 33]. In addition to these metrics, we evaluate the novel T-mAP metric, which addresses the shortcomings of previous metrics, as discussed in Section 3.1.
F.3 Backbones
We use three types of architectures in our experiments. The first is the GRU network [5], one of the top-performing neural architectures for event sequences [1]. We also implement the continuous-time LSTM (CT-LSTM) from the NHP method [19]. Since CT-LSTM requires a specialized loss function and increases training time, we use it exclusively with the NHP method, preferring GRU in other cases. An additional advantage of GRU is that its output is equal to its hidden state, simplifying the estimation of intermediate hidden states for autoregressive inference starting from the middle of a sequence. Lastly, we implement the AttNHP continuous-time transformer model [36].
F.4 Datasets